# Forest-of-Thought: Scaling Test-Time Compute for Enhancing LLM Reasoning

### v1

@@ 5 Conclusion
This paper introduces a novel method, the Forest of Thought (FoT), aimed at significantly enhancing the reasoning capabilities of large language models (LLMs). FoT leverages a structured framework that integrates multi-path exploration and dynamic activation of reasoning paths, addressing key limitations in existing LLM reasoning paradigms. This enables the model to achieve robust and efficient problem-solving across complex tasks while generating diverse reasoning outcomes without relying on backpropagation or fine-tuning.
By integrating multiple independent “thinking” models, such as the Tree of Thought and Monte Carlo Trees, FoT effectively tackles complex problems and incorporates a sparse activation mechanism to substantially improve computational efficiency. Additionally, this paper systematically investigates the scaling relationship between reasoning time and accuracy in LLMs, providing a theoretical foundation for optimizing their reasoning performance.
@@ Appendix B Example Prompts - GSM8K and MATH
Tabel 6 demonstrate a clear trend: as the number of activated subtrees grows, the error rate decreases, reflecting the enhanced reasoning capability and robustness of the FoT approach. The scaling law comparison across various methods underscores the efficiency of FoT in utilizing additional computational resources, particularly when contrasted with other models that exhibit diminishing returns or slower error reduction as complexity increases. This suggests that the FoT framework scales more effectively with increased subtree activation, making it a powerful tool for addressing challenging reasoning tasks like those in GSM8K.

### v2

@@ 5 Conclusion
This paper introduces Forest of Thought (FoT), a novel framework designed to significantly enhance the reasoning capabilities of large language models (LLMs). FoT employs a structured framework that integrates multi-path exploration and dynamic activation of reasoning paths, addressing key limitations in existing LLM reasoning paradigms. This approach enables models to achieve robust and efficient problem-solving across complex tasks while generating diverse reasoning outcomes—without relying on backpropagation or fine-tuning.
By incorporating reasoning frameworks such as Tree of Thought and Monte Carlo Tree Search, FoT effectively tackles complex problems while leveraging a sparse activation mechanism to significantly improve computational efficiency. Furthermore, this paper systematically explores the scaling relationship between reasoning time and accuracy in LLMs, providing a theoretical foundation for optimizing their reasoning performance.
Impact Statement
The FoT framework introduces a novel and highly efficient approach to improving the reasoning capabilities of LLMs. By leveraging multiple reasoning trees, dynamic self-correction strategies, and consensus-based decision-making, FoT enables LLMs to solve complex problems with higher precision and reduced computational cost. The key innovation lies in the adaptive reasoning process that allows the model to explore multiple paths, detect and correct errors in real-time, and efficiently narrow down the search space. This approach has significant implications for fields that require complex decision-making, such as mathematics, logic, and AI-driven problem solving, leading to more robust and reliable AI systems.
@@ Appendix B Example Prompts - GSM8K and MATH
Tabel 7 demonstrate a clear trend: as the number of activated subtrees grows, the error rate decreases, reflecting the enhanced reasoning capability and robustness of the FoT approach. The scaling law comparison across various methods underscores the efficiency of FoT in utilizing additional computational resources, particularly when contrasted with other models that exhibit diminishing returns or slower error reduction as complexity increases. This suggests that the FoT framework scales more effectively with increased subtree activation, making it a powerful tool for addressing challenging reasoning tasks like those in GSM8K.